4) Your firm is a new retail space in it. Each busi ness will pay a fixed strip mall and you have multiple businesses that would like lease each month. The cash flows from each of the businesses has approximately the same amounn of names, square footnge requirements, and monthly expected cash flows for esch ot tive amount of rent each month plus a percentage of the gross sales businesses that would like to lease space in your strip Square Feet Required Expected Monthly Cash Business Name Videos Now Flow 70,000 52.500 4000 500 2500 1500 500180.000 12,000 Pizza Warehouse 28,500 30 1/2 Flavors S-Mart Wal Verde Drugs 6000 14.000 1000 Multigular Wireless If your new strip mall will have 16,000 square feet of should you lease and why? retail space available to be leased, to which businesses a) Is your combination of leases most efficient, and why? (Hint: Does your current combination best exhaust the space? Is there any other combination yields to higher cash flow? b)
Answer #1

Solution:

Cash flow per square feet

A)

70000/4000 = 17.5
52500/3500 = 15
52500/2500 = 21
25500/1500 = 17
28500/1500 = 19
180000/12000 = 15
147000/6000 = 24.5
22250/1000 = 22.25

We will lease first with the highest cash flow per square feet
Hence, we will lease the following

WalVerde Drugs for 6000 square feet

Multigular Wireless for 1000 square feet

Pizza Warehouse for 2500 square feet

30 ½ Flavors for 1500 square feet

Videos Now for 4000 square feet.

Unused space is 1000 square feet

B)

The above combination is very efficient because when we are leasing the space, our objective should be to maximize the cash flows from activities I.e, from rentals aswell as a percentage and also from sales.

The other combinations do not yield to higher cash flows as compared to the above combination mentioned in the answer (a), so the answer is no.